OpenStudy (anonymous):

Water freezes at thirty-two degrees Fahrenheit, or zero degrees Celsius. Use temperature conversion formulas to show that these two measures are indeed the same temperature

OpenStudy (anonymous):

help me

OpenStudy (campbell_st):

well a simple conversion formula is \[C = \frac{5}{9} \times (F -32)\]

OpenStudy (anonymous):

I tried the formula it did not work

OpenStudy (anonymous):

I need help solving

OpenStudy (campbell_st):

wow... so \[C = \frac{5}{9} \times (32 - 32)\] doesn't give 0

OpenStudy (campbell_st):

I think you need to check your working... and order of operation

OpenStudy (anonymous):

help